OpenStudy (anonymous):

What is an equation of a parabola with the given vertex and focus? vertex: (–2, 5); focus: (–2, 6)

OpenStudy (anonymous):

First you wanna use: \[ x^2 = 4py \]Where \(p\) is the distance between the focus and the vertex.

OpenStudy (anonymous):

his will give you a parabola that is centered at the origin.

OpenStudy (anonymous):

Then you want to translate it by replacing \(y\) with \((y-k)\) and \(x\) with \((x-h)\), where the vertex is \((h,k)\)

OpenStudy (anonymous):

then what?? @wio

OpenStudy (anonymous):

Do all that and you're left with a parabola.

OpenStudy (anonymous):

what equation do i use to find it out?

OpenStudy (anonymous):

\(p\) is the distance from the vertex and focus, in this case it is \(1\). So we start out with \[ x^2=4(1)y = 4y \implies y= \frac{x^2}{4} \]

OpenStudy (anonymous):

Then we translate our parabola into the correct vertex. \[ y - 5 = \frac{(x-(-2))^2}{4} = \frac{(x+2)^2}{4}\implies y = \frac{1}{4}(x+2)^2+5 \]Then foil

OpenStudy (anonymous):

soo would that be the answer? @wio